# Exploiting the Uncoordinated Privacy Protections of Eye Tracking and VR Motion Data for Unauthorized User Identification

### v1

@@ 6 Discussion
First, we establish that both VR motion data and eye tracking data are feasible modalities for user identification, and that combining data streams could improve biometric authentication performance when one data stream is more informative than the other. These findings are consistent with prior work [14] and demonstrate the potentially complementary nature of VR motion data and eye tracking data for user identification. We then show that existing privacy mechanisms [33, 23] can provide demonstrable privacy protections when they are applied to every available data stream. However, failing to provide comprehensive privacy in this way can lead to the novel vulnerability that we explore in this work, where the unprotected data streams are used to circumvent the privacy protections afforded by the protected data streams. The level of biometric authentication performance for this exploitation is comparable to the performance observed when no privacy mechanisms were applied at all, which emphasize the potential ramifications that this exploitation has on user privacy. In a real attack scenario, an adversary would not need to know which data streams were unprotected, or if any of the data streams are protected at all. As long as at least one informative data stream is unprotected (or insufficiently protected, in the case of our hand data), they would still be able to identify users with above-chance accuracy. By emphasizing these implications of our findings, we hope to motivate the development of privacy solutions that are robust to the presence of auxiliary data, especially when that auxiliary data is correlated, synchronized, and as informative as the target data stream.
We observe that EKYT tends to produce better biometric performance when gaze data is provided to the model, but also has more difficulty extracting identifying information from privatized gaze streams than any other data. We note that EKYT was originally created and validated for biometric authentication for gaze data, so there may be some artifacts of the architecture design or data processing that may complicate comparisons between its performance on gaze data nd non-gaze data. However, we note that we still observed relatively high biometric performance on VR motion data, even when we did not extensively fine-tune the model or optimally pre-process the head and hand data to achieve the highest possible biometric performance. While our data processing strategy was sufficient for the current investigation, it is possible that biometric performance could additionally be improved by applying sensor fusion techniques to more explicitly define relationships between the data streams. Pre-processing the input data to explicitly identify user-specific information shared across data streams may help the model learn these correlations more effectively, making the combined data streams more powerful than simply the sum of their individual contributions—this is an additional consideration for user privacy that must be addressed in future work.
@@
We utilized two existing privacy mechanisms that were validated in prior work—in particular, these mechanisms both produced favorable privacy-utility trade-offs and demonstrated viability for real-time use in their original works. We observed that the mechanism we applied to VR data was less effective at degrading EKYT’s biometric performance. Nair et al.’s [23] privacy mechanism for VR motion data was designed to obfuscate select physical characteristics from the VR motion data, which potentially leaves identifying behavioral patterns available to leverage for user identification. Although obscuring a user’s height and wingspan showed some demonstrable privacy-enhancing effect, EKYT was able to extract other identifiable patterns from the data, potentially from the behavioral patterns that were left unprotected. This represents an opportunity for advancement in the field of privacy mechanisms for streaming VR data—future work in this direction may explore privacy mechanisms that can obfuscate both physical characteristics and behavioral patterns from VR motion data.
@@
Ideally, a comprehensive privacy solution for VR would involve multiple layers of privacy protection, so that no single point of failure may enable privacy violations. In addition to applying privacy mechanisms to obfuscate identifiable data and implementing system-level privacy protections mentioned in Section 2, a VR platform may further restrict which applications may access these data streams based on their stated purpose. Applications that do not demonstrate the need for high-granularity, sample-to-sample streaming data can be provided abstractions of the data stream, which contain less sensitive information that could possibly be exploited. This strategy has been proposed separately for both gaze data [4] and gesture-based motion data [7], and can be combined with other existing techniques to help address the emerging privacy concerns that we highlight in this work. At the same time, we emphasize the need for privacy solutions that are intentionally designed to address the unique privacy considerations of multi-sensor environments, so that privacy protections can be formalized, guaranteed, and consistently delivered, rather than simply happen by coincidence.
@@ 7 Limitations and Future Directions
Although we demonstrate that combining data streams in VR can improve biometric performance when only a portion of data streams are privatized, it is unclear whether it can also allow the model to resolve biometric claims with higher levels of confidence. Reducing a biometric model’s confidence below its decision threshold can be another promising avenue to preventing unauthorized user identification. To further investigate whether these data streams contain complementary information for biometrics, future work may consider reporting measures of confidence alongside the biometric performance of their models. Future work in this direction can also improve the explainability of their results by either using statistical, feature-based biometrics models or utilizing explainable AI techniques [1] to provide more insight for deep learning-based biometric models.
Due to the limited availability of data sets that contain gaze, headset and hand data in VR, the insights gained in this investigation is limited by the relatively small number of subjects used for biometric authentication. Our investigation utilized a larger number of subjects than most of the prior work in literature—compared to our corpus of 38 users, similar studies perform biometric authentication on populations comprising 12 [14], 15 [24], 19 [27], and 43 [32] people. However, we recognize that having access to a data set with more users can enable a more robust investigation of user identification in VR. This work can be expanded on by utilizing larger data sets that contain more users who are recorded across multiple sessions, even if these data sets only feature only two of the three data streams that we explored in this work. We also limited the scope of this investigation to explore a novel privacy consideration in VR environments because of the limited availability of data, but the insights from this study can reasonably be applied to analogous situations in augmented or mixed reality.
@@ 8 Conclusion
We explored a novel vulnerability in VR systems that stems from a patchwork approach to developing privacy-enhancing techniques for the data that is collected on these platforms. Namely, when privacy protections are applied to only a subset of data streams available on the platform, users can still be vulnerable to identification attacks when unprotected sensor data is used to circumvent them. The complementary nature of eye tracking, headset tracking, and hand tracking data allows adversaries to bypass targeted privacy measures and infer user identity with accuracy levels comparable to a scenario where no privacy mechanism is applied at all. These findings highlight the need for comprehensive privacy-enhancing techniques that account for the broad range of potentially correlated data streams collected in VR. In doing so, future work can advance the state-of-the-art in user privacy for VR ecosystems and contribute to the continued effort to grant users autonomy over their own data in these environments.

### v3

@@ 6 Discussion
We first establish that both VR motion data and eye tracking data are suitable for both user verification and identification in our problem setting, and that combining data streams can improve biometric authentication performance when one data stream is more informative than the other. These findings are consistent with prior work [19] and demonstrate the potentially complementary nature of VR motion data and eye tracking data for user identification.
@@
We then show that existing privacy mechanisms from [41, 31] can provide demonstrable privacy protections when they are applied to each available data stream. However, failing to apply these privacy mechanisms to one or more data streams can lead to privacy leakage, which can enable an adversary to circumvent the privacy protections afforded by the protected data streams by leveraging the sensitive information in the unprotected data streams. In addition, privacy mechanisms that do not impart significant privacy protection (as observed with our hand data) can create a similar vulnerability. We observed that the level of biometric performance when such privacy leakage is present is comparable to the performance observed when no privacy mechanisms were applied at all, which emphasize the potential ramifications that this vulnerability has on user privacy. In a real attack scenario, an adversary would not need to know which data streams were unprotected, or if any of the data streams are protected at all. By emphasizing these implications of our findings, we hope to motivate the development of privacy solutions that are robust to the presence of auxiliary data, especially when that auxiliary data is correlated, synchronized, and complementary in nature.
While this work demonstrates the effect that cross-sensor privacy leakage has on the risk of unauthorized user re-identification in VR platforms, we note that the results of this investigation are heavily influenced by our selection of our privacy mechanisms, authentication model, and data pre-processing strategy, and that some of the trends we observe here may not generalize to other combinations of the three.
@@
First, we utilized two existing privacy mechanisms whose privacy-utility trade-offs and viability for real-time use were validated in prior work. We observed that the mechanism we applied to VR data was less effective at degrading EKYT’s biometric performance overall. Nair et al.’s [31] privacy mechanism for VR motion data was designed to obfuscate select physical characteristics, which potentially leaves identifying behavioral patterns available to leverage for user identification. Although obscuring a user’s height and wingspan showed some demonstrable privacy-enhancing effect, EKYT was able to extract other identifiable patterns from the data, potentially from the behavioral patterns that were left unprotected. This represents an opportunity for advancement in the field of privacy mechanisms for streaming VR data—future work in this direction may explore privacy mechanisms that can obfuscate both physical characteristics and behavioral patterns from VR motion data.
Secondly, we observe that EKYT tends to produce higher biometric performance when gaze data is provided to the model, but also experiences the most degradation in biometric performance when using privatized gaze streams than any other data. EKYT was originally created and validated for biometric authentication for gaze data, so there may be some artifacts of the architecture design or data processing that may complicate comparisons between its performance on gaze data and non-gaze data. Although we did not extensively fine-tune our pre-processing of VR motion data to achieve the highest possible biometric performance, we observed relatively high biometric performance on VR motion data. However, it is possible that biometric performance could additionally be improved by applying sensor fusion techniques to more explicitly define relationships between the data streams. Pre-processing the input data to explicitly identify user-specific information shared across data streams may help the model learn these correlations more effectively, making the combined data streams more powerful than simply the sum of their individual contributions.
Finally, and most importantly, we note in Section 5.3 that the volume of data that was privatized affected EKYT’s ability to effectively learn to extract user-specific behaviors. Namely, when all eleven input channels are available and all nine channels of VR motion data are privatized, EKYT cannot effectively extract user-specific information from the remaining two channels of unmodified gaze data (E20). Conversely, when only the two channels of gaze data are privatized, the amount of privacy leakage represented by the nine unmodified channels of VR motion data produce significantly better biometric performance (E17). While it seems straightforward that privatizing more data generally results in lower levels of privacy leakage, we explicitly note that this is the direct result of the way we provided this input data to EKYT. If we revised our methodology to enforce that each test condition contains the same number of privatized and unmodified data streams, we anticipate that we would observe less variation in the experiments evaluating the effect of incomplete privacy protection. This phenomenon highlights one of the many complexities of preserving user privacy in multi-sensor environments like VR. Although each individual sensor may only contribute a small amount of potentially identifiable information, the sheer volume and variety of data that is collected about a user through these sensors can collectively create an unexpectedly detailed profile of a user over time. These sensors may perform separate functions to enable different applications in VR, but they can capture potentially complementary information about behavioral idiosyncrasies that may then manifest as privacy vulnerabilities that are difficult to anticipate.
@@
Ideally, a comprehensive privacy solution for VR would involve multiple forms of privacy protection, so that there is no single point of failure that may enable privacy violations. In addition to applying privacy mechanisms to obfuscate identifiable data and implementing system-level privacy protections mentioned in Section 2, a VR platform can restrict which applications may access these data streams based on their stated purpose. Applications that do not demonstrate need for sample-to-sample streaming data can be provided abstractions of the data stream which contain less sensitive information that could possibly be exploited. This strategy has been proposed separately for both gaze data [5] and gesture-based motion data [8], and can be combined with other existing techniques to help address the emerging privacy concerns that we highlight in this work. Ultimately, our work emphasizes the need for privacy solutions that intentionally address the unique privacy considerations of multi-sensor environments, so that privacy protections can be formalized, guaranteed, and experienced consistently.
@@ 7 Limitations and Future Directions
Although we demonstrate that combining data streams in VR can improve biometric performance when only a portion of data streams are privatized, it is unclear whether it can also allow the model to resolve biometric claims with higher levels of confidence. Reducing a biometric model’s confidence below its decision threshold can be another promising avenue to preventing unauthorized user identification. To further investigate whether these data streams contain complementary information for biometrics, future work should report measures of confidence alongside the biometric performance of their models. Future work in this direction can also improve the explainability of their results by either using statistical, feature-based biometrics models or utilizing explainable AI techniques [1] to provide more insight for deep learning-based biometric models.
Due to the limited availability of data sets that contain both gaze data and VR motion data, this investigation has the limitation of only reporting biometric performance for a relatively small number of subjects. Compared to our corpus of 38 users, similar studies perform biometric authentication on populations comprising 12 [19], 15 [32], 19 [35], and 43 [40] people. Although our investigation features one of the largest sample sizes available, we acknowledge that having access to a data set with more users can enable a more robust investigation of user identification in VR. This work can be expanded by utilizing larger data sets that contain more users who are ideally recorded across multiple sessions, even if these data sets only feature only two of the three data streams that we explored in this work. We also limited the scope of this investigation to explore a novel privacy consideration in VR environments because of this limitation, but the insights from this study can reasonably be applied to analogous situations in augmented or mixed reality.
Finally, the results presented in this work may have limited generalizability, as they were produced using a single biometric authentication model working on a single data set, and evaluated only one privacy mechanism per data stream. While it is clear from our results that the privacy leakage we observe directly impact user authentication rates in this experiment, the extent to which this relationship is observed may vary significantly depending on the authentication model employed, the nature and parameters of the privacy mechanisms that are applied, and any auxiliary pre-processing that a data set may undergo in addition to privacy enhancement. Future work in this direction may involve the development and validation of robust, generalizable methods to formally estimate the amount of risk users face as a result of privacy leakage, to enable the development of straightforward “best practices” for preserving user privacy applicable to a wider variety of data sets and sensor types.
@@ 8 Conclusion
We conducted an empirical study exploring a vulnerability in VR systems that stems from the incomplete application of privacy-enhancing techniques to data that is collected on these platforms. Namely, when privacy protections are applied to only a subset of data streams available on the platform, users can still be vulnerable to identification attacks when unprotected sensor data is used to circumvent them. We demonstrate that the complementary nature of eye tracking, headset tracking, and hand tracking data allows adversaries to bypass targeted privacy measures and infer user identity with accuracy levels comparable to a scenario where no privacy mechanism is applied at all. These findings highlight the need for comprehensive privacy-enhancing techniques that account for the collection of synchronized, potentially correlated data streams VR platforms. In doing so, future work can advance the state-of-the-art in user privacy for VR ecosystems and contribute to the continued effort to grant users autonomy over their own data in these platforms.